1. Now suppose the car doubles its speed to 25.0 m/s (~90 km/hr). Its brakes can apply a force of 5.00x103 N.  What is the minimum stopping distance at this speed?
1
Expert's answer
2021-05-18T11:09:29-0400

Apply Newton's second law to determine deceleration:


"a=\frac Fm."

The minimum stopping distance at this speed is


"d=\frac{v^2}{2a}=\frac{v^2m}{2F}=0.125m,"

where m is the mass of the car.